### v1

# Relaying Quantum Information

Abstract
Quantum relays are central to both quantum communication and distributed quantum computing, enabling long-distance transmission and modular architectures. Unlike classical repeaters, quantum repeaters preserve coherence without amplifying quantum information, relying on entanglement swapping and quantum error correction to overcome loss and decoherence. In this work, we investigate the transmission of quantum information via quantum relay channels. Our three-terminal relay model captures the trade-off between repeater-assisted and repeaterless communication strategies. We propose a decode-forward coding scheme and analyze both entanglement-assisted and unassisted scenarios. Our framework allows for different entanglement topologies between the transmitter, the relay and the destination receiver, recovering known results on entanglement-assisted and unassisted communication. Furthermore, we discuss the interpretation of coding with quantum side information. These findings serve as a stepping stone for the design of secure, efficient, and reliable quantum networks and the practical realization of quantum repeaters and long-range quantum key distribution.
I Introduction
Relaying quantum information plays a central role in quantum communication and distributed quantum computing, enabling long-distance transmission and modular architectures [1, 2]. Quantum information technology faces significant challenges due to the fragile nature of quantum states, which are highly susceptible to noise. In quantum computing, fault tolerance requires robust quantum error correction (QEC) codes and efficient syndrome extraction methods [3]. However, practical challenges such as correlated errors and slow decoding limit current QEC performance [4]. Distributed relays could facilitate high-rate quantum error correction by enabling concurrent decoding across separate quantum devices [5], yet demand coherent and reliable quantum state transfer across them [6]. Moreover, reliable long-range quantum communication is severely limited by photon loss and environmental noise, in both fiber-optic and free-space transmission [7]. By leveraging entanglement swapping [8] and teleportation-based protocols [9], quantum relays not only help maintain coherence and entanglement in distributed architectures, but also extend the distance between end users [10, 11, 12].
From a security standpoint, a the US National Security Agency (NSA) has highlighted several challenges regarding the current applicability of quantum cryptography [13]. Among these challenges is the need for trusted intermediate stations in current implementations of quantum key distribution (QKD). As pointed out by Renner and Wolf  [14], this will change once quantum repeaters are employed. These devices work entirely on the quantum level and hence do not need to be trusted. Therefore, using quantum repeaters will prevent security risks from insider threats. Understanding the relaying of quantum information over quantum channels can serve as a significant milestone on the road to developing quantum repeaters.
In order to distribute entanglement and send quantum information, the quantum relay would need to operate as a quantum repeater [15]. A quantum repeater is a device that enables long-distance quantum communication by dividing the transmission into shorter segments and using entanglement swapping and purification to extend quantum correlations without directly amplifying quantum information, which is fundamentally prohibited by the no-cloning theorem. Unlike classical repeaters, they do not amplify signals but instead rely on QEC or entanglement-based techniques to overcome loss and decoherence [16, 17]. A basic comparison between the classical and quantum repeaters is illustrated in Fig. 1.
Figure 1: Comparison of classical and quantum repeater architectures. Classical repeaters amplify signals, while quantum repeaters utilize entanglement.
Cooperation in quantum communication networks has become a major focus of research in recent years, driven by advances in both experimental techniques and theoretical developments [18, 19, 20]. Entanglement serves as a valuable resource for network communication [21]. In the point-to-point setting, entanglement assistance between the transmitter and receiver can significantly boost throughput [22], even when the entanglement resource is noisy [23] or unreliable [24].
In multihop networks, multiple rounds of communication take place, making synchronization a crucial requirement [25, 26]. Consider a central clock that ticks $n$ times. Between clock ticks $i-1$ and $i$, the sender and relay transmit their respective channel inputs, $A_{i}$ and $D_{i}$ (see Fig. 2). At tick $i$, the relay and destination receive the channel outputs, $E_{i}$ and $B_{i}$, respectively. A short delay is thus necessary to ensure causal operation. The classical relay channel model was originally introduced by van der Meulen [27], and later adopted as a foundation for multihop network models [28].
Causality constraints in network settings often require the use of block Markov coding [29], where the transmitter sends a sequence of blocks, and each block encodes information related to both current and previous transmissions. Quantum versions of block Markov coding have been studied in the context of classical-quantum (c-q) relay channels [30, 31], quantum communication with parameter estimation [32], and quantum cribbing scenarios [33].
An information-theoretic framework for primitive quantum relay channels was first introduced in [34], and later extended in [35] to address classical communication over fully quantum relay channels using various coding strategies. Other quantum repeater models can be found in [36, 37, 38, 39, 40].
In general, the quantum communication literature can broadly be categorized into two approaches: repeater-aided and repeaterless transmission. The Pirandola–Laurenza–Ottaviani–Banchi (PLOB) bound [41] sets the fundamental rate-distance limit for repeaterless communication, i.e., without a quantum repeater. The repeater-aided transmission enables the distribution of quantum correlations across long distances, by dividing the channel into shorter segments and placing where intermediate relay terminals. Through entanglement swapping, the quantum repeaters generate entanglement and maintain coherence without violating the no-cloning theorem [42, 43].
Experimental realizations for repeaterless and repeater-aided communication were demonstrated in Refs.  [44] and Refs.  [45], respectively.
In this work, we focus on the transmission of quantum information over quantum relay channels, aiming to characterize the trade-offs between repeater-assisted and repeaterless communication scenarios. For the discrete-variable case, we study the three-terminal quantum relay channel illustrated in Fig. 3. This type of relay network serves as a fundamental building block in cooperative communication  [46] and can also be used to establish entanglement between network nodes  [47]. Specifically, we consider the decode-forward strategy, in which the relay fully decodes the quantum message received from Alice, re-encodes it, and forwards it to Bob. We analyze both the entanglement-assisted and unassisted cases. For the unassisted case, we leverage the fact that communication occurs over multiple rounds, allowing the protocol to reuse entanglement generated as a by-product in block $i-1$ during block $i$.
Our model also permits entanglement between the inputs of Alice and the relay (see Fig. 2), which as we show leads to improved achievable transmission rates over the case of the orthogonal links (see Ref. [36] for an example of a repeater chain composed of orthogonal links). As a special case, our results recover those of Dupuis in Ref. [48], who studied the capacity of quantum channels with side information at the transmitter. From the perspective of quantum error correction, our setup can also be interpreted such that the undisturbed reference system $R$ corresponds to the quantum data register, and the ancillary system $M$, transmitted through the communication channel, enables subsequent error syndrome extraction.
The paper is organized as follows. In Sec. II, we present preliminary definitions and the quantum relay channel model. Sec. III describes the coding framework for strictly causal quantum relaying. Our main results for the decode-forward strategy, including assisted and unassisted scenarios, are given in Sec. IV. Sec. V provides an overview of the fully quantum Slepian-Wolf (FQSW) protocol, a central tool in our analysis. Detailed proofs of the achievable rates appear in Sec. VI and Sec. VII. Finally, Sec. VIII provides a summary and concluding remarks.
Figure 2: Coding for a quantum relay channel $\mathcal{N}_{AD\rightarrow BE}$.
II Preliminaries and Notation
In this section, we introduce the basic definitions and notation used throughout the paper. A quantum state is represented by a density operator $\rho$ acting on a Hilbert space $\mathcal{H}$, satisfying $\operatorname{Tr}(\rho)=1$ and $\rho\succeq 0$. We assume all Hilbert spaces to be finite-dimensional. The set of all such density operators is denoted by $\mathcal{D}(\mathcal{H})$.
The outcome distribution of a quantum measurement can be described using a positive operator-valued measure (POVM), which is a collection of positive semi-definite operators $\{\Delta_{j}\}$ such that $\sum_{j}\Delta_{j}=\mathbbm{1}$, where $\mathbbm{1}$ is the identity operator. According to the Born rule, if the system is in the state $\rho$, the probability of observing outcome $j$ is given by $\Pr(j)=\operatorname{Tr}(\Delta_{j}\rho).$. The quantum entropy of a density operator $\rho$ is defined as $H(\rho)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}-\operatorname{Tr}(\rho\log\rho),$. which reduces to the Shannon entropy of the eigenvalue distribution of $\rho$. For a quantum state $\rho_{AB}$ defined over the joint system $A$ and $B$ with $\rho_{AB}\in\mathcal{D}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$, the quantum mutual information is defined as
$I(A;B)_{\rho}=H(\rho_{A})+H(\rho_{B})-H(\rho_{AB}),$
where $\rho_{A}=\operatorname{Tr}_{B}(\rho_{AB})$ and $\rho_{B}=\operatorname{Tr}_{A}(\rho_{AB})$ are the reduced density operators. The conditional quantum entropy is given by $H(A|B)_{\rho}=H(\rho_{AB})-H(\rho_{B})$, and the coherent information is defined as $I(A\rangle B)_{\rho}=H(\rho_{B})-H(\rho_{AB})_{\rho}=-H(A|B)_{\rho}$.
We use the following notation for state approximation: We write $\rho\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon}}}{{\approx}}}\sigma$ iff ${\|\rho-\sigma\|}\leq\varepsilon$. Uhlmann’s theorem is a useful tool in this work: If the reduced states of ${|{\varphi}\rangle}_{AB}$ and ${|{\psi}\rangle}_{AC}$ satisfy $\varphi_{A}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon}}}{{\approx}}}\psi_{A}$, then there exists a partial isometry $V_{C\rightarrow B}$ such that
${|{\varphi}\rangle}_{AB}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle 2\sqrt{\varepsilon}}}}{{\approx}}}V_{C\rightarrow B}{|{\psi}\rangle}_{AC}.$
In this work, we consider a fully quantum relay channel as a model for the three-terminal network depicted in Fig. 3. A quantum relay channel is described by a completely positive trace-preserving (CPTP) map $\mathcal{N}_{AD\rightarrow BE}$, where $A$, $D$, $B$, and $E$ correspond to the sender’s transmitter, the relay’s transmitter, the destination receiver, and the relay’s receiver, respectively (see Fig. 3). We assume that the channel is memoryless. Specifically, if the sender and relay inputs are given by the systems $A^{n}=(A_{1},A_{2},\ldots,A_{n})$ and $D^{n}=(D_{1},D_{2},\ldots,D_{n})$ across $n$ channel uses, then the joint input state $\rho_{A^{n}D^{n}}$ evolves under the tensor product channel $\mathcal{N}_{A^{n}D^{n}\rightarrow B^{n}E^{n}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\left(\mathcal{N}_{AD\rightarrow BE}\right)^{\otimes n}$. The specific communication setting will be formalized in Sec. III, where we assume that the relay encodes information in a strictly causal manner. That is, at each time step $i$, the relay first transmits $D_{i}$ and only then receives $E_{i}$.
III Coding Scheme
Consider a quantum relay channel $\mathcal{N}_{AD\rightarrow BE}$. We define a code for the transmission of quantum information via $\mathcal{N}_{AD\rightarrow BE}$, from Alice to Bob. A $(2^{nQ},n)$ quantum code for $\mathcal{N}_{AD\rightarrow BE}$ is defined by:
  • Finite Hilbert space $\mathcal{H}_{M}$ of dimension $\dim(\mathcal{H}_{M})=2^{nQ}$.
  • An encoding channel $\mathcal{E}_{M\rightarrow A^{n}}:\mathcal{D}(\mathcal{H}_{M})\rightarrow\mathcal{D}(\mathcal{H}_{A}^{\otimes n})$, which encodes the state of the information system $M$ into the input system $A^{n}$.
  • A sequence of strictly casual encoding maps $\Gamma^{(i)}_{E_{i-1}\bar{E}^{i-2}\rightarrow{\bar{E}^{i-1}D_{i}}}:\mathcal{D}(\mathcal{H}^{\otimes(i-1)}_{E})\rightarrow\mathcal{D}(\mathcal{H}^{\otimes(i-1)}_{E}\otimes\mathcal{H}_{D})$, with $\bar{E}^{i-1}=(E_{1},...,E_{i-1})$ representing all quantum states up to communication round $i$.
  • A decoding channel $\mathcal{P}_{B^{n}\to\hat{M}}$ acting on $\mathcal{H}_{B}^{\otimes n}$.
We denote the code by $\mathscr{C}=(\mathcal{E},\Gamma,\mathcal{P})$.
The communication scheme is illustrated in Fig. 2. The systems $A^{n}$ and $D^{n}$ represent the transmissions from the sender and the relay, respectively, while $E^{n}$ and $B^{n}$ denote the received outputs at the relay and the receiver. The system $R$ represents the complementary system, to which the state in the relay communication scheme may be entangled. The joint state at the beginning of the protocol is described by the pure state ${|{\Psi}\rangle}\in\mathcal{D}(\mathcal{H}_{M}\otimes\mathcal{H}_{R})$. The system Alice, representing the sender, selects a quantum system $\rho_{M}$ which is intended for Bob, representing the destination receiver. The encoding is done by Alice as $\rho_{M}\rightarrow\mathcal{E}_{M\rightarrow A^{n}}(\rho_{M})$, followed by the transmission of the $A^{n}$ systems over $n$ channel uses.
The transmission is done iteratively, at step $i$, the relay encodes the sequence of states $(E_{i-1},\bar{E}^{i-2})$ using strictly casual encoding map $\Gamma^{(i)}_{E_{i-1}\bar{E}^{i-2}\rightarrow{\bar{E}^{i-1}D_{i}}}$, and then receives $E_{i}$ for $1\leq i\leq n$. Each step $i$ is composed of two operations, the action of $\Gamma^{(i)}$ and the action of the quantum channel $\mathcal{N}$.
The first time step $i=1$ requires an "initial condition" for the encoding map $\Gamma^{(1)}$ as:
$\rho_{RA^{n}D_{1}}=\left[\left(\mathrm{id}_{R}\otimes\mathcal{E}_{M\rightarrow A^{n}}\right)(\Psi_{RM})\right]\otimes\Gamma_{E_{0}\rightarrow D_{1}}^{(1)}(1),$ (1)
where $\Psi\equiv{|{\Psi}\rangle\!\,\langle{\Psi}|}$ and $E_{0}$ is a degenerate system of dimension 1.
In what follows, we employ the notation $A^{n}=A^{q-1}_{1}\bigcup A^{n}_{q}$, where $A^{i}_{j}=\{A_{p}\}^{i}_{p=j}$. For example, taking $q=2$, we split $A^{n}$ into $A_{1}$ and $A_{2}^{n}\equiv\{A_{p}\}_{p=2}^{n}$. At $i=1$, the relay receives $E_{1}$ with the state, after the action of the quantum channel:
$\rho_{RB_{1}E_{1}A_{2}^{n}}=\left[\mathrm{id}_{R}\otimes\mathcal{N}_{AD\rightarrow BE}\otimes\mathrm{id}_{A_{2}^{n}}\right]\rho_{RA_{1}D_{1}A_{2}^{n}}.$ (2)
The second time step $i=2$ begins with the action of $\Gamma^{(2)}$, corresponding to the relay encoding of $D_{2}$ as:
$\rho_{R\bar{E}_{1}D_{2}A_{2}^{n}B_{1}}=\left[\mathrm{id}_{R}\otimes\Gamma_{E_{1}\rightarrow\bar{E}_{1}D_{2}}^{(2)}\otimes\mathrm{id}_{A_{2}^{n}B_{1}}\right]\rho_{RB_{1}E_{1}A_{2}^{n}}.$ (3)
The action of the quantum channel $\mathcal{N}$ at time step $i=2$ is done by sending the state $E_{2}$ to the relay as:
$\rho_{RB_{2}E_{2}A_{3}^{n}B_{1}\bar{E}_{1}}=\\ \left[\mathrm{id}_{R}\otimes\mathcal{N}_{AD\rightarrow BE}\otimes\mathrm{id}_{A_{3}^{n}B_{1}\bar{E}_{1}}\right]\rho_{RA_{2}D_{2}A_{3}^{n}B_{1}\bar{E}_{1}}.$ (4)
At the final time step $i=n$, the relay encodes $D_{n}$ as:
$\rho_{R\bar{E}^{n-1}D_{n}A_{n}^{n}B^{n-1}}=\\ \left[\mathrm{id}_{R}\otimes\Gamma_{E_{n-1}\bar{E}^{n-2}\rightarrow\bar{E}^{n-1}D_{n}}^{(n)}\otimes\mathrm{id}_{A_{n}B^{n-1}}\right]\\ (\rho_{RE_{n-1}\bar{E}^{n-2}A_{n}B^{n-1}}),$ (5)
then the state $E_{n}$ is sent to the relay as the last step of the communication protocol:
$\rho_{RB^{n}E_{n}\bar{E}^{n-1}}=\\ \left[\mathrm{id}_{R}\otimes\mathcal{N}_{AD\rightarrow BE}\otimes\mathrm{id}_{B^{n-1}\hat{G}_{E}^{n-1}}\right]\rho_{RA_{n}D_{n}B^{n-1}\hat{G}_{E}^{n-1}}.$ (6)
The output state for Bob (destination receiver) is the reduced state $\rho_{B^{n}}=\operatorname{Tr}_{RE_{n}\bar{E}^{n-1}}\left[\rho_{RB^{n}E_{n}\bar{E}^{n-1}}\right]$. Bob receives the channel output state $\rho_{B^{n}}$ over the output system $B^{n}$, applies the decoding channel $\mathcal{P}_{B^{n}\to\hat{M}}$ and recovers $\hat{M}$ as the estimation of the quantum system sent by Alice. The error of the code is:
$e^{(n)}(\Psi,\mathscr{C})=\frac{1}{2}||\Psi_{RM}-\Psi_{R\hat{M}}||_{1}.$ (7)
IV Main Results
Here we present our main results for the quantum relay channel $\mathcal{N}_{AD\rightarrow BE}$. We establish achievable rates that are based on different coding strategies, with and without the entanglement assistance.
Figure 3: Three-terminal quantum relay schematic. The transmitters at the sender and relay are denoted by $A$ and $D$, respectively, while the receivers at the relay and destination are labeled $E$ and $B$.
IV-A Rate-Limited Entanglement Assistance
In the decode-forward case, the relay decodes the entire information. Consider the three party communication scheme shown in Fig. 3. Let $\mathcal{N}_{AD\rightarrow BE}$ be a quantum channel, and let ${|{\sigma}\rangle}_{AA_{1}D}$ be any pure state with $\omega_{A_{1}BE}=\mathcal{N}_{AD\rightarrow BE}(\sigma_{A_{1}AD})$.
Proposition 1.
Assume that Bob has access to entanglement assistance represented by entanglement consumption and generation rates $L_{B}$ and $\hat{L}_{B}$, respectively. Then any rate point $(Q,L_{B},\hat{L}_{B})$ such that:
$\begin{split}Q\leq H(A_{1}|D)_{\sigma}-L_{B}+\hat{L}_{B}\\ Q\leq I(A_{1}\rangle E)_{\omega}-L_{B}-\hat{L}_{B}\\ Q\leq I(A_{1}\rangle B)_{\omega}+L_{B}-\hat{L}_{B},\end{split}$ (8)
is achievable for transmission with rate-limited entanglement assistance, where $A_{1}$ is an arbitrary auxiliary system.
The first condition states that the quantum information transmitted by Alice, along with the entanglement consumed from her share of the EPR pairs, must fit within the channel input available to Alice, conditioned on the relay’s input to the channel. Specifically, let $\Delta L_{B}=L_{B}-\hat{L}_{B}$ denote the net entanglement consumption. Then, the first condition states that $Q+\Delta L_{B}\leq H(A_{1}|D)_{\sigma}$. The third condition says that the channel can transmit $Q-\Delta L_{B}\leq I(A_{1}\rangle B)$ qubits per channel use, when taking the cost of entanglement into account, hence subtracting the net entanglement consumption. The second condition directly reflects the bottleneck introduced by the relay, where both $L_{B}$ and $\hat{L}_{B}$ appear with the negative sign. This arises due to the nature of our Block Markov coding scheme, where the entanglement assistance available to Bob in block $j$ is effectively generated by the relay in block $j-1$. See Sec. VI-B.
IV-B Without Entanglement Assistance
Our main result is given below. Define
$\mathsf{Q}_{\text{D-F}}(\mathcal{N})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max_{\sigma}\left[\min\left\{I(A_{1}\rangle B)_{\omega},I(A_{1}\rangle E)_{\omega}\right\}\right],$
where the maximum is taken over all pure states of the form ${|{\sigma_{A_{1}AD}}\rangle}$, and where $\omega_{A_{1}BE}=\mathcal{N}_{AD\rightarrow BE}(\sigma_{A_{1}AD})$. The subscript ‘D-F’ stands for decode-forward bound.
Theorem 2.
The quantum capacity of the quantum relay channel $\mathcal{N}_{AD\rightarrow BE}$ without assistance satisfies:
$C(\mathcal{N})\geq\mathsf{Q}_{\text{D-F}}(\mathcal{N}).$ (9)
IV-C Unlimited Entanglement Assistance
In the next theorem, we consider communication of quantum information via the quantum relay channel, when Alice and Bob are provided with pre-shared entanglement assistance, in which case, $L_{B}>\hat{L}_{B}>0$ (see Proposition 1). Let
$\mathsf{Q}_{\text{EA,D-F}}(\mathcal{N})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max_{\sigma}\left[\frac{1}{2}I(A_{1};B)_{\omega}-\frac{1}{2}I(A_{1};D)_{\sigma}\right],$
where the maximization is as before, and the subscript ‘EA’ indicates the availability of entanglement assistance.
Theorem 3.
The entanglement-assisted quantum capacity of the quantum relay channel $\mathcal{N}_{AD\rightarrow BE}$ satisfies:
$C_{\text{EA}}(\mathcal{N})\geq\mathsf{Q}_{\text{EA,D-F}}(\mathcal{N}).$ (10)
Remark 1.
The theorem above has an immediate consequence on the transmission of classical information with entanglement assistance. Specifically, using the superdense coding protocol, classical information can be sent at a rate of $R\geq$ $2\cdot\mathsf{Q}_{\text{EA,D-F}}(\mathcal{N})$ (classical information bits per transmission).
Remark 2.
The input state ${|{\sigma}\rangle}_{AA_{1}D}$ represents the correlation that is generated between the transmitter and the relay. Intuitively, we think of the auxiliary system $A_{1}$ as Alice’s entanglement resource. Due to entanglement monogamy, if the resource $A_{1}$ is maximally entangled with the relay’s transmission, $D$, then there is no correlation with the channel input, $A$. Similarly, if the resource is maximally entangled with the channel input, then it has no correlation with the relay. Therefore, it is not immediately obvious how to distribute the correlation between $A_{1}$, $A$, and $D$.
Taking ${|{\sigma}\rangle}_{AA_{1}D}={|{\phi_{A_{1}A}}\rangle}\otimes{|{\psi}\rangle}_{D}$ yields the direct transmission bound:
$\mathsf{Q}_{\text{D-F}}(\mathcal{N})\geq\max_{\sigma}\left[I(A_{1}\rangle B)_{\omega}\right]$
i.e., the coherent information of the channel from Alice to Bob. Similarly, we recover the entanglement-assisted-capacity:
$\mathsf{Q}_{\text{EA,D-F}}(\mathcal{N})\geq\max_{\sigma}\left[I(A_{1};B)_{\omega}\right].$
On the other hand, ${|{\sigma}\rangle}_{AA_{1}D}={|{\psi}\rangle}_{A}\otimes{|{\xi_{A_{1}D}}\rangle}$ results in a zero rate, as $H(A_{1}|D)$ in Proposition 1 becomes non-positive.
Remark 3.
In our communication setting (see Fig. 3), the links are not orthogonal; that is, $\mathcal{N}_{AD\rightarrow BE}\neq\mathcal{P}_{A\rightarrow E}\otimes\mathcal{M}_{D\rightarrow B}$. In contrast, if the channel links were orthogonal, i.e., $\mathcal{N}_{AD\rightarrow BE}=\mathcal{P}_{A\rightarrow E}\otimes\mathcal{M}_{D\rightarrow B}$, Theorem 2 yields:
$\mathsf{Q}_{\text{D-F}}(\mathcal{N})\geq\min\left\{I_{c}(\mathcal{M}),I_{c}(\mathcal{P})\right\},$
in the unassisted case. The orthogonal links model was previously considered by Ghalaii and Pirandola [36, 38], Gyongyosi and Imre [37, 39], and Shi et al. [40]. Intuitively, the generalized communication model allows for entanglement between the inputs of Alice and the relay, allowing for a higher achievable rate.
Remark 4.
The results presented in Theorem 2 and Theorem 3 bear resemblance to the entanglement-assisted quantum capacity of a quantum channel with side information at the transmitter, as studied in Ref. [48]. In the side information model, the channel $\tilde{\mathcal{N}}_{DA\to B}$ is governed by an external environment $D$, which is entangled with the transmitter. The connection becomes apparent if we interpret the transmission to the relay as entangling the transmitter with the channel’s environment (see Fig. 3). Note, however, that the environment state in the side information model [48] is fixed, i.e., $\sigma_{D}$ is dictated by the model. In contrast, our formula involves optimization as we are free to choose the entanglement constellation between the transmitter and the relay.
In the upcoming sections, we provide a detailed proof for the achievable rates in Theorem 2 and Theorem 3.
V Analytic Tools: The FQSW Theorem
Before presenting the achievability proof, we provide a quick overview of the fully quantum Slepian-Wolf (FQSW) protocol as presented in Ref. [49]. This summary broadly follows the one given in Ref. [48]. Consider a pure state ${|{\psi}\rangle}_{ABC}$ shared by Alice, Bob, and a reference system held by Charlie. Now, Alice wishes to send her state to Bob with high fidelity and use as few qubits as possible. In the FQSW protocol, Alice first applies a unitary $U_{A\rightarrow A_{1}A_{2}}$, drawn according to the Haar measure, and then sends $A_{2}$ to Bob.
The FQSW result is a decoupling theorem. The protocol induces a state such that the $A_{1}$ and $C$ sub-systems are “almost" in a product state at the end of the protocol. As shown in Ref. [49], one may evaluate how close the resulting state is to a decoupled state by:
$\mathbb{E}_{U}\left\|\rho_{A_{1}C}-\frac{\mathbbm{1}_{A_{1}}}{|A_{1}|}\otimes\psi_{C}\right\|^{2}_{1}\leq\frac{|A||C|}{|A_{2}|^{2}}\operatorname{Tr}\left[\left(\psi_{AC}\right)^{2}\right]$ (11)
where the expectation is taken over the random unitary $U_{A\rightarrow A_{1}A_{2}}$, according to the Haar measure. Applying this protocol to a block of $n$ transmissions from Alice to Bob, we have $\operatorname{Tr}\left[\left(\psi_{AC}\right)^{2}\right]\approx 2^{-nH(AC)_{\psi}}$ for sufficiently large $n$, by the Quantum AEP Property [50].
Since Eq. (11) holds on average, there must exist at least one unitary that satisfies the bound. That is, using the state-approximation notation, there exists $U_{A\rightarrow A_{1}A_{2}}$ such that
$\rho_{A_{1}C}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon}}}{{\approx}}}\frac{\mathbbm{1}_{A_{1}}}{|A_{1}|}\otimes\psi_{C}$ (12)
with the proximity parameter $\varepsilon=\frac{|A||C|}{|A_{2}|^{2}}\operatorname{Tr}\left[\left(\psi_{AC}\right)^{2}\right]$.
Since Bob holds the purifying system of $A_{1}C$, there exists a local unitary $U^{\prime}$ that splits his purifying system into two separate purifying systems of the two sub-systems $A_{2}$ and $C$, respectively. The purifying system of $C$ is exactly the system that Alice wanted to send, $A_{2}$, together with Bob’s system $B$. Whereas, $A_{1}$ becomes half of a maximally entangled pair, shared by Alice and Bob. Therefore, as a by-product of the FQSW protocol, we obtain an entanglement resource shared by Alice and Bob at the end of the protocol.
VI Proof of Proposition 1
First, we introduce the notation for the entanglement resources consumed and generated during the protocol. The quantum systems $G_{E}$ and $G_{B}$ represent the pre-shared entanglement resources available at the relay and the destination receiver (Bob), respectively, with associated consumption rates defined as $L_{E}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{n}\log|G_{E}|$ and $L_{B}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{n}\log|G_{B}|$.
In accordance with Sec. V, the quantum systems $\hat{G}_{1,A}$ and $\hat{G}_{2,A}$ represent the by-product entanglement generated between the sender (Alice) and the relay, and between the sender and the destination receiver, respectively. These systems are located at the sender, with corresponding generation rates given by $\hat{L}_{E}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{n}\log|\hat{G}_{1,A}|$ and $\hat{L}_{B}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{n}\log|\hat{G}_{2,A}|$.
VI-A Decoupling Properties
We begin with decoupling properties. Let ${|{\sigma_{A_{1}AD}}\rangle}$.
VI-A1 Decoupling at the relay’s transmitter
Consider the product state, ${|{\sigma_{A_{1}AD}}\rangle}^{\otimes n}$, where $A_{1}$ is the encoder ancilla, $A$ the channel input, and $D$ the relay output. Projecting onto the typical subspace of $A_{1}$ gives:
${|{\psi_{A_{1,\text{typ}}^{n}A^{n}D^{n}}}\rangle}=(\Pi_{A_{1}^{n}\to A_{1,\text{typ}}^{n}}\otimes\mathbbm{1}_{AD}){|{\sigma_{A_{1}AD}}\rangle}^{\otimes n}.$ (13)
As in Sec. V, we apply a random unitary $U_{A_{1,\text{typ}}^{n}\to RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}}$, where $R$ is the reference system with $|R|=2^{nQ}$. The resulting state is
${|{\theta_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}A^{n}D^{n}}}\rangle}=\left(U\otimes\mathbbm{1}_{A^{n}D^{n}}\right){|{\psi_{A_{1,\text{typ}}^{n}A^{n}D^{n}}}\rangle}.$ (14)
Considering the reduced state $\theta_{RG_{E}G_{B}D^{n}}$, we can “decouple" between $D^{n}$ and $RG_{E}G_{B}$, and write:
$\theta_{RG_{E}G_{B}D^{n}}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{1}}}}{{\approx}}}\frac{\mathbbm{1}_{RG_{E}G_{B}}}{|RG_{E}G_{B}|}\otimes\sigma_{D^{n}},$ (15)
which, by the FQSW Theorem (see Sec. V), holds for
$\varepsilon_{1}\equiv\varepsilon_{1}(n)$ $=\frac{|A_{1,\text{typ}}^{n}|\cdot|D^{n}|}{|\hat{G}_{1,A}|^{2}|\hat{G}_{2,A}|^{2}}\cdot 2^{-nH(A_{1}D)}$
$\leq 2^{-n(H(A_{1}|D)_{\sigma}-L_{E}-L_{B}+\hat{L}_{E}+\hat{L}_{B}-Q-\delta)},$ (16)
and sufficiently large $n$. The last inequality follows from standard typical projector properties (see Property 15.1.2 in Ref. [51]). Observe that $\varepsilon_{1}(n)$ tends to zero as $n\to\infty$, provided that
$Q<H(A_{1}|D)_{\sigma}-L_{E}-L_{B}+\hat{L}_{E}+\hat{L}_{B}-\delta.$ (17)
VI-A2 Decoupling at the relay’s receiver
Consider a Stinespring representation of the quantum relay channel $\mathcal{N}_{AD\to BE}$, with an isometry $V^{\mathcal{N}}_{AD\to BEJ_{O}}$. Applying this isometry produces the following output state:
${|{\omega_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}B^{n}E^{n}J_{O}^{n}}}\rangle}=\\ (\mathbbm{1}\otimes\left(V^{\mathcal{N}}_{AD\to BEJ_{O}})^{\otimes n}\right){|{\theta_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}A^{n}D^{n}}}\rangle},$ (18)
where $R$ is Alice’s resource, $E^{n}$ is received at the relay, $B^{n}$ at the destination receiver, and $J_{O}^{n}$ represents the environment of the joint output, $B^{n}E^{n}$. We denote the composite environment of the relay by $J_{E}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}BJ_{O}$, and consider the reduced state $\phi_{R\hat{G}_{1,A}\hat{G}_{2,A}G_{B}J_{E}^{n}}$. Decoupling the relay’s environment:
$\phi_{R\hat{G}_{1,A}\hat{G}_{2,A}G_{B}J_{E}^{n}}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{2}}}}{{\approx}}}\frac{\mathbbm{1}_{R\hat{G}_{1,A}\hat{G}_{2,A}G_{B}}}{|R\hat{G}_{1,A}\hat{G}_{2,A}G_{B}|}\otimes(\mathcal{N}_{A\to J_{E}}({|{\sigma_{A}}\rangle}))^{\otimes n},$ (19)
where
$\varepsilon_{2}(n)$ $=\frac{|A_{1,typ}||J_{E}^{n}|}{|G_{E}|^{2}}\cdot 2^{-nH(A_{1,typ}J_{E})}$
$\leq 2^{-n(I(A_{1}\rangle E)_{\omega}+L_{E}-L_{B}-\hat{L}_{E}-\hat{L}_{B}-Q-\delta)},$ (20)
for sufficiently large $n$, where the inequality follows from the same arguments as used in Sec. VI-A1. The last expression tends to zero as $n\to\infty$, if
$Q<I(A_{1}\rangle E)_{\omega}+L_{E}-L_{B}-\hat{L}_{E}-\hat{L}_{B},$ (21)
with respect to $\omega_{A_{1}BE}=\mathcal{N}_{AD\rightarrow BE}(\sigma_{A_{1}AD})$.
VI-A3 Decoupling at the destination receiver
We now consider the overall transmission from Alice to Bob, treating the relay’s output as part of Bob’s environment. The output state is as in Eq. (18). Denote Bob’s environment by $J_{B}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}EJ_{O}$. Similarly as before, decoupling Bob’s environment yields
$\phi_{R\hat{G}_{1,A}\hat{G}_{2,A}G_{E}J_{B}^{n}}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{3}}}}{{\approx}}}\frac{\mathbbm{1}_{R\hat{G}_{1,A}\hat{G}_{2,A}G_{E}}}{|R\hat{G}_{A}G_{E}|}\otimes(\mathcal{N}_{A\to J_{B}}({|{\sigma_{A}}\rangle}))^{\otimes n},$ (22)
where
$\varepsilon_{3}(n)$ $=\frac{|A_{1,\mathrm{typ}}||J_{B}^{n}|}{|G_{B}|^{2}}\cdot 2^{-nH(A_{1,\mathrm{typ}}J_{B})}$
$\leq 2^{-n(I(A_{1}\rangle B)_{\omega}-L_{E}+L_{B}-\hat{L}_{E}-\hat{L}_{B}-Q-\delta)},$ (23)
for sufficiently large $n$. This tends to zero if
$Q<I(A_{1}\rangle B)_{\omega}-L_{E}+L_{B}-\hat{L}_{E}-\hat{L}_{B}.$ (24)
VI-B Encoding and decoding procedures
In this section, we describe the encoding and decoding procedures for the sender, relay, and destination receiver. We use a Block Markov scheme: Alice transmits a sequence of $T$ blocks $A^{n}(1),\ldots,A^{n}(T)$ each block is of size $n$, in order to distribute $M(2),\ldots,M(T)$, each of dimension $2^{nQ}$. The effective transmission rate is thus $\frac{T-1}{T}Q$, which tends to $Q$ in the limit of $T\to\infty$.
Throughout, we use the following convention for block indexing to clearly distinguish between the current and past instances of a quantum system during the $T$-block encoding and decoding process. A quantum system labeled as $S$ refers to the system at block index $j$, i.e., $S\stackrel{{\scriptstyle\mathrm{def}}}{{=}}S(j)$. A quantum system labeled as $S^{\prime}$ refers to the system at the previous block, i.e., $S^{\prime}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}S(j-1)$.
VI-B1 Transmitter, encoder and channel isometries
For each block $j$, there exists a family of partial isometries $W_{MG_{A}\rightarrow A^{n}\hat{G}_{1,A}\hat{G}_{2,A}}$ such that, according to the Sec. VI-A1, the following approximation holds:
${|{\theta_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}A^{n}D^{n}}}\rangle}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{1}}}}{{\approx}}}\\ \left(W_{MG_{1,A}G_{2,A}\rightarrow A^{n}\hat{G}_{1,A}\hat{G}_{2,A}}\otimes\mathbbm{1}\right)\cdot\\ \left(\Phi_{RM}\otimes\Phi_{G_{1,A}G_{E}}\otimes\Phi_{G_{2,A}G_{B}}\otimes\sigma_{D^{n}}\right),$ (25)
where $\Phi_{RM}$ is the maximally entangled message state between the reference $R$ and the “message" register $M$, and $\Phi_{G_{1,A}G_{E}}\otimes\Phi_{G_{2,A}G_{B}}$ is the entanglement assistance shared between Alice and the relay, and between Alice and Bob. The output systems $A^{n}\hat{G}_{1,A}\hat{G}_{2,A}$ jointly purify the reduced state $\theta_{RG_{E}G_{B}D^{n}}$.
The channel output is then
${|{\omega_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}B^{n}E^{n}J_{O}^{n}}}\rangle}=\\ \left[(V^{\mathcal{N}}_{AD\to BEJ_{O}})^{\otimes n}\otimes\mathbbm{1}\right]{|{\theta_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}A^{n}D^{n}}}\rangle}.$ (26)
VI-B2 Relay encoder and decoder
We begin with the relay encoder. For each block $j-1$, there exists a family of partial isometries $\Lambda_{\tilde{M}^{\prime}\hat{G}_{E}^{\prime n}\rightarrow{D}^{n}G_{E}^{\prime}}$ such that the following approximation holds:
${|{\theta_{R^{\prime}G_{E}^{\prime}G_{B}^{\prime}\hat{G}_{1,A}^{\prime}\hat{G}_{2,A}^{\prime}A^{\prime n}{D}^{n}}}\rangle}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{2}}}}{{\approx}}}\left[\Lambda_{\tilde{M}^{\prime}\hat{G}_{E}^{\prime n}\rightarrow{D}^{n}G_{E}^{\prime}}\otimes\mathbbm{1}\right]\\ \cdot\Phi_{R^{\prime}\tilde{M}^{\prime}}\otimes\Phi_{\hat{G}_{1,A}^{\prime}\hat{G}_{E}^{\prime n}}\otimes\Phi_{\hat{G}_{2,A}^{\prime}\hat{G}_{B}^{\prime n}}\otimes\psi_{B^{\prime n}J_{O}^{\prime n}},$ (27)
where $\theta_{D^{n}}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{1}}}}{{\approx}}}\sigma_{D}^{\otimes n}$, based on Sec. VI-A2. This is important for the next stage.
As for the relay decoder, invoking the results of Sec. VI-A2, the partial isometries $\Gamma_{E^{n}G_{E}\rightarrow\tilde{M}\hat{G}_{E}^{n}}$ satisfy
$\Phi_{R\tilde{M}}\otimes\Phi_{\hat{G}_{1,A}\hat{G}_{E}^{n}}\otimes\psi_{\hat{G}_{2,A}G_{B}B^{n}J_{O}^{n}}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{2}}}}{{\approx}}}\\ \left[\Gamma_{E^{n}G_{E}\rightarrow\tilde{M}\hat{G}_{E}^{n}}\otimes\mathbbm{1}\right]\cdot{|{\omega_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}B^{n}E^{n}J_{O}^{n}}}\rangle},$ (28)
where the quantum system $\tilde{M}$ is the quantum “message restoration" system of the relay.
VI-B3 Destination decoder
Similarly to the above, according to Sec. VI-A3, for each block $j$, there exists a family of partial isometries $\Delta_{B^{n}G_{B}\rightarrow\hat{M}\hat{G}_{B}^{n}}$ such that:
$\Phi_{R\hat{M}}\otimes\Phi_{\hat{G}_{2,A}\hat{G}_{B}^{n}}\otimes\psi_{\hat{G}_{1,A}G_{E}E^{n}J_{O}^{n}}\mathrel{\stackrel{{\scriptstyle\mathclap{\textstyle\varepsilon_{3}}}}{{\approx}}}\\ \left[\Delta_{B^{n}G_{B}\rightarrow\hat{M}\hat{G}_{B}^{n}}\otimes\mathbbm{1}\right]\cdot{|{\omega_{RG_{E}G_{B}\hat{G}_{1,A}\hat{G}_{2,A}B^{n}E^{n}J_{O}^{n}}}\rangle},$ (29)
where the quantum system $\hat{M}$ is the quantum “message estimation" system of the destination receiver, producing $\approx\Phi_{R\hat{M}}$ between Alice and Bob. ∎
VII Proof of Theorem 2 and Theorem 3
In this section we prove Theorem 2 and Theorem 3 based on our result in Proposition 1.
VII-A Unassisted case
Taking the entanglement assistance rates to be $L_{E}=\hat{L}_{E}\rightarrow 0$ and $L_{B}=\hat{L}_{B}\rightarrow 0$, and combining the constraints in Eq. (17), Eq. (21), and Eq. (24) yields the following bound:
$Q<\min\left\{I(A_{1}\rangle B)_{\omega},I(A_{1}\rangle E)_{\omega}\right\}-\delta.$ (30)
The inequality in Eq. (17) is inactive, since $H(A_{1}|D)_{\sigma}=I(A_{1}\rangle A)_{\sigma}$ by entropy duality, and $I(A_{1}\rangle A)_{\sigma}\geq I(A_{1}\rangle B)_{\omega}$ by the data processing inequality. This concludes the proof of Theorem 2.
VII-B Unlimited assistance
Assuming that only Bob has access to entanglement assistance, we set the relay’s entanglement rates to be $L_{E}=\hat{L}_{E}\rightarrow 0$. Isolating $Q$ in Eq. (24) and Eq. (17), we achieve
$Q<\frac{1}{2}I(A_{1};B)_{\omega}-\frac{1}{2}I(A_{1};D)_{\sigma}-\delta.$ (31)
This concludes the proof. ∎
VIII Summary and Discussion
We study quantum information transmission over a three-terminal quantum relay network, focusing on the decode-forward coding strategy in both entanglement-assisted and unassisted settings. Our model captures key tradeoffs between repeater-assisted and repeaterless communication, and highlights the role of entanglement between the transmitter and the relay. Furthermore, our framework generalizes previous results, which focus on either repeaterless transmission or repeater-aided communication. These insights serve as a stepping stone towards the design of quantum networks (in the long-term, a quantum Internet), and provide a foundation for practical quantum repeaters and long-distance QKD.
One avenue for future research is to develop alternative coding strategies. In particular, we may require the relay to decode only a part of the information. This is particularly relevant when the direct channel from Alice to Bob is not strictly noisier (e.g., degraded) than the channel to the relay. Classical information theory suggests that there is often an optimal balance between (a) the amount of information decoded through the relay and (b) the amount decoded independently by the destination receiver [52, 35]. Exploring how this trade-off manifests in the quantum setting may yield new insights into quantum information transmission over relay channels.
Acknowledgments
Y. Ilin and U. Pereg were supported by ISF, n. 939/23, OMC, n. 86160946, DIP, n. 2032991, and Helen Diller Quantum Center, n. 2033613. U. Pereg was also supported by Chaya Chair, n. 8776026, and VATAT Program for Quantum Science and Technology, n. 86636903.